A college sports foundation deposited a total of $24,000 into two simple interest accounts. The annual simple interest rate on one account is 7%. The annual simple interest rate on the second account is 11%. How much is invested in each account if the total annual interest earned is 19% of the total investment?
Answer by mananth(16949)

19% of 24000 = 4560
Part I 7.00% per annum ------------- Amount invested =x
Part II 11.00% per annum ------------ Amount invested = y
24000
Interest----- 4560

Part I 7.00% per annum ---x *
Part II 11.00% per annum ---y
Total investment
x + 1 y= 24000 -------------1
Interest on both investments
7.00% x + 11.00% y= 4560
Multiply by 100
7 x + 11 y= 456000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -168000.00
Add this to (2)
0 x 4 y= 288000
divide by 4
y = 72000 - x)
Part I 7.00% $ -48000
Part II 11.00% $ 72000

CHECK
-48000 --------- 7.00% ------- -3360.00
72000 ------------- 11.00% ------- 7920.00
Total -------------------- 4560.00

m.ananth@hotmail.ca